Ernest spent a total of $64 for a pair of jeans and a shirt. The jeans cost $6 more than the shirt. What was the cost of the jeans?: Ernest spent a total of $64 for a pair of jeans and a shirt. The jeans cost $6 more than the shirt. What was the cost of the jeans?
Answer by mangopeeler07(442)

s=price of shirt
j=price of jeans

s+j=64------------------------------Ernest spent a total of $64 for a pair of jeans and a shirt.

j=s+6-------------------------------The jeans cost $6 more than the shirt.

Plug in s+6 for j in the first equation:
s+j=64
becomes
s+s+6=64

Combine like terms
2s+6=64

Subtract 6 from both sides
2s=58

Divide both sides by 2
s=29

s=price of shirt
s+6=price of jeans

Answer:
shirt=$29
*jeans=$35*